The of a platinum resistance Thermometer at o degree is 8.4 ohm,at 100 degree is 10 ohm and 16.5 at the temperature of liquid inside a bath.find the temperature of liquid.the thermometer is then placed in a flask containing liquid hydrogen at -150 degree.what will be the value of its resistance?.


1
Expert's answer
2021-04-05T11:07:47-0400

"\alpha=\frac{\Delta R}{R_0 \Delta t}=\frac{10-8.4}{8.4\cdot (100-0)}=1.9\cdot10^{-3}~°C^{-1},"


"R_1=R_0(1+\alpha t_1)\implies t_1=\frac{R_1-R_0}{\alpha R_0},"

"t_1=\frac{16.5-8.4}{1.9\cdot 10^{-3}\cdot 8.4}=508°C,"


"R_2=R_0(1+\alpha t_2),"

"R_2=8.4\cdot (1+1.9\cdot10^{-3}\cdot (-150))=6~\Omega."